Integration by Parts Fractions


1.) ∫4dx/(x^4-1)


2.) ∫8dx/x(x^2+2)^2


1
Expert's answer
2021-05-07T09:10:56-0400

1) ∫ 4 dx / (x^4 - 1)

We know that x4 - 1 = (x2 -1) * (x2 +1)

Also (x2 +1) - (x2 -1) = 2

So the above integral can be written as


2 ∫ [(x2+1) - (x2-1)] dx / (x2+1) * (x2-1)

2 ∫ [1/(x2-1) - 1/(x2+1) ] dx

2 ln(x-1/x+1) - 2 arctan (x) + c [Since ∫ dx/x2 - a2 = (1/2a ) ln(x-a/ x+a) + c

and ∫ dx/a2+x2 = (1/a ) arctan(x/a) + c ]



2) ∫ 8 dx/ x(x2+2)2

The above integral can be expressed as a sum of partial fractions as shown below

8 / x(x2+2)2 = A/x + Bx/(x2+2) + Cx/ (x2+2)2

On comparing the coefficients in LHS and RHS and solving them we have

A = 2

B = -2

C = -4

Now the above integral takes the form

∫ [ 2/x - 2x/(x2+2) - 4x/ (x2+2)2 ] dx

Let I1 = ∫ 2 dx/x

I2 = ∫ 2x dx/(x2+2)

I3 = ∫ 4x dx/(x2+2)2 ................(1)



I1 = 2 ln(x)



For I2 ,

Let x2+2 = t

2x dx = dt

∴ I2 = ∫ dt / t

= ln(t)

= ln( x2+2 ) ....................(2)



For I3 ,

Let x2+2 = u

2x dx = du


∴ I3 = 2 ∫ du / u2

= - 2 /u

= -2/ (x2+2) ................(3)




∴ ∫ 8 dx/ x(x2+2)2 = I1 - I2 - I3

= 2 ln(x) - ln (x2 + 2) + 2 / (x2 + 2) + c

= ln[ x2 / (x2 + 2)] + 2 / (x2+2) + c